if leah is 6 years older than Sue and JOhn is 5 years older then leah, and the total of their ages is 41. How old is Sue?
Can someone clarify this for me?Thank you

Answer by rfer(16322)

Set them all to start from the same person.
Sue=x=8
Leah=x+6=14
John=x+11=19
x+(x+6)+(x+11)=41
3x+17=41
3x=41-17
3x=24
x=8